OpenStudy (anonymous):

Check answer 4 medal A function is shown below: f(x) = (0.07)x What does the function represent? Exponential growth of 7% Exponential decay of 7% <---- Exponential growth of 93% Exponential decay of 93%

OpenStudy (anonymous):

or D?

OpenStudy (anonymous):

I think were keeping the 0.07 so that must mean 93% is decaying

OpenStudy (anonymous):

Definitely A.

OpenStudy (anonymous):

i thinkj

OpenStudy (anonymous):

its not increasing cause the base is less than 1

OpenStudy (anonymous):

it is A for sure

OpenStudy (igreen):

@awesomeness812 @briannamooney is correct, it's decreasing, not increasing.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I thought so , why are people picking A

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

whats the next question

OpenStudy (anonymous):

btw I was asking if it was A. Cz I have that same question on my paper

OpenStudy (anonymous):

anything you need help with anything else

OpenStudy (igreen):

I’m going with D. 1 here will be the same thing as 0. 1 means it doesn’t increase or decrease. A 7% decrease would be subtracting 0.07 from 1: f(x) = (1 – 0.07)^x f(x) = (0.93)^x That’s the function for a 7% decrease. However, a 93% decrease would be subtracting 0.93 from 1: f(x) = (1 – 0.93)^x f(x) = (0.07)^x That’s the function for a 93% decrease, so D it is.

OpenStudy (igreen):

@briannamooney

OpenStudy (anonymous):

..

OpenStudy (anonymous):

@iGreen .-.-

OpenStudy (igreen):

What? @awesomeness812

OpenStudy (igreen):

You do know your answer was COMPLETELY wrong, right?

OpenStudy (igreen):

@briannamooney

OpenStudy (anonymous):

Sorry was afk cause i went to get my notes and yeah im going with D also , thank you !

OpenStudy (igreen):

No problem!

OpenStudy (anonymous):

nvm